# Almost Surely Safe Alignment of Large Language Models at Inference-Time

### v2

@@ D.1 Experiment Setup
We conducted all experiments on NVIDIA V100 GPUs, utilizing Python 3.11.11 and CUDA 12.4, for model development and training. The experimental setup was based on the PKU-SaferRLHF dataset.Our experiments involved two primary models: the unsafe-aligned reproduced Alpaca-7B model [^6] https://huggingface.co/PKU-Alignment/alpaca-7b-reproduced (Taori et al., 2023) and the safe-aligned Beaver-v3-7B model (Ji et al., 2024b). We employed the reward model Khanov et al. (2024) [^7] https://huggingface.co/argsearch/llama-7b-rm-float32 and a cost model Ji et al. (2024b) [^8] https://github.com/PKU-Alignment/safe-rlhf during the training stage for critic network, test-time inference and evaluation stages.
D.2 Baselines and Hyper-parameter Settings
The baseline methods we compare against include BoN, Beam Search, RECONTROL (Kong et al., 2024), and ARGS (Khanov et al., 2024). For BoN, we use a strategy where the top N outputs are sampled, and the best result is selected based on a predefined criterion. Similarly, Beam Search explores multiple beams during the search process and selects the best output based on a beam-width parameter. In RECONTROL, an MLP network is employed as the value network to intervene in the decision-making process, guiding the generation through reinforcement learning (Kong et al., 2024). ARGS, on the other hand, implements a logits-based greedy token-wise search strategy, where tokens are generated sequentially based on the maximum likelihood of the next token (Khanov et al., 2024).
Given the limited research on safe alignment with inference-time methods, we adapt these baseline methods to enable safe inference, ensuring a fair comparison. To this end, we incorporate a Lagrangian multiplier term to control the balance between reward and cost, allowing for safe inference based on their open-source implementations. Notably, BoN and Beam Search utilize a form of blocking sampling, while ARGS and RECONTROL employ token sampling methods.
In our setup, we modify the inference process of the baseline methods with a Lagrangian multiplier by using the following score for token selection: $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ where $\lambda$ is the Lagrangian multiplier, that controls the influence of the safety cost score $\mathcal{C}_{\text{safety}}$. We unify $\lambda=5$ and also the sampling parameters for baseline methods and InferenceGuard to ensure a fair comparison. Each method has its own specific settings, as follows:
@@ D.2 Baselines and Hyper-parameter Settings
ARGS 1 20 N/A
RECONTROL 1 N/A n steps = 30, step size = 0.5
BoN 32 100, 200, 500 N/A
Beam Search 32 128, 256 D=128, K=$N/4$
InferenceGuard 32 128, 512 M=2, D=128, K=$N/4$
@@
These hyperparameter settings ensure a fair and controlled comparison of the baseline methods and the proposed InferenceGuard model. We compare the performance of baseline methods, with reward model, cost model and lagriangian multiper setting, under these hyperparameter settings, and summarised the result in Table 2.
When sampling from the base model, we used a temperature of 1.0 without top_k nor top_p.
Table 2: Performance Comparison on Dataset PKU-SafeRLHF

### v3

@@ D.1 Experiment Setup
All the experiments are conducted on our internal cluster using Python 3.11 and bfloat16 mixed-precision training. The setup used a single high-memory compute device, and the inference process can be reproduced on any similar hardware with at least 64 GB of on-device memory available. We assess safe inference across three datasets that vary in their safety sensitivity: 1) PKU-SafeRLHF with safety-critical instructions for reward and cost alignment, 2) HEx-PHI: mixed-sensitivity health and philosophy instructions with challenging edge cases, and 3) HH-RLHF with high-safety prompts with strong human preferences. Our experiments involved four major language models with varying safety alignment levels: the unaligned Alpaca-7B model [^5] https://huggingface.co/PKU-Alignment/alpaca-7b-reproduced[67], the safety-aligned Beaver-v3-7B model[70] [^6] https://huggingface.co/PKU-Alignment/beaver-7b-v3.0, the unaligned Vicuna-7B[68] [^7] https://huggingface.co/lmsys/vicuna-7b-v1.5 and Llama3.1-8B-Instruct models [69] [^8] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. We employed two sets of reward models: 1) Llama2-based reward model llama-7b-rm[43] [^9] https://huggingface.co/argsearch/llama-7b-rm-float32 and cost model beaver-7b-unified-cost [70] [^10] https://huggingface.co/PKU-Alignment/beaver-7b-unified-cost, and 2) Llama-3-based reward model and cost model QRM-LLama-3.1-8B[71] [^11] https://huggingface.co/nicolinho/QRM-Llama3.1-8B during the training stage for critic network, test-time inference and evaluation stages.
D.2 Experiment Settings and Results
The baseline methods we compare against include BoN, Beam Search, RE-Control[42], and ARGS[43]. For BoN, we use a strategy where the top N outputs are sampled, and the best result is selected based on a predefined criterion. Similarly, Beam Search explores multiple beams during the search process and selects the best output based on a beam-width parameter. In RE-Control, an MLP network is employed as the value network to intervene in the decision-making process, guiding the generation through reinforcement learning[42]. ARGS, on the other hand, implements a logits-based greedy token-wise search strategy, where tokens are generated sequentially based on the maximum likelihood of the next token[43].
Given the limited research on safe alignment with inference-time methods, we adapt these baseline methods to enable safe inference using the Lagrangian and safety augmentation approaches, ensuring a fair comparison. To this end, we incorporate a Lagrangian multiplier term or a safety augmentation based on their open-source implementations to enable safety-aligned inference. Notably, BoN and Beam Search utilize a form of blocking sampling, while ARGS and RE-Control employ token sampling methods.
In the Lagrangian-based method setup, we modify the inference process of the baseline methods by incorporating a Lagrangian multiplier, using the following score function: $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ where $\lambda$ is the Lagrangian multiplier, that controls the influence of the safety cost score $\mathcal{C}_{\text{safety}}$. For the main experiment results, we set $\lambda=5$ uniformly and fix the sampling parameters across all baseline methods and InferenceGuard to ensure a fair comparison. Each method also has additional specific settings, as detailed below:
@@ D.2 Experiment Settings and Results
ARGS 1 128 N/A
RE-Control 1 N/A n steps = 30, step size = 0.5
BoN 32 64, 128, 256 N/A
Beam Search 32 64, 128, 256 D=128, K=$N/4$
InferenceGuard 32 64, 128, 512 M=2, D=128, K=$N/4$
@@
We compare the performance of baseline methods with lagrangian-based methods and safety augmentation methods, under these hyperparameter settings, and summarize the result in Table 2 for Alpaca-7B and Beaver-7B, Table 3 for Vicuna-7B, and Table 4 for LLaMA-3.1-8B-Instruct. When sampling from the base model, we used a temperature of 1.0 without top_k nor top_p.
Table 2: Performance Comparison using Alpaca-7B and Beaver-7B, evaluated by reward models llama-7b-rm and beaver-7b-unified-cost, on Dataset PKU-SafeRLHF $N=128$, $\lambda=5$
Method Average Reward Safety Rate Inference Time (s)
Alpaca-7B Base 6.15 ($\pm$ 1.51) 29.47% 0.7
RE-Control 6.2 ($\pm$ 1.56) 29.5% 1.13
RE-Control + Lagrangian multiplier 6.19 ($\pm$ 1.50) 29.7% 1.33
Best-of-N + Lagrangian multiplier 5.31 ($\pm$ 1.62) 49.37% 8.0
Best-of-N + Augmented safety 7.22 ($\pm$ 1.90) 58.57% 9.75
Beam search + Lagrangian multiplier 6.58 ($\pm$ 1.95) 50.19% 14.5
Beam search + Augmented safety 8.29 ($\pm$ 2.02) 58.89% 14.6
ARGS $\omega=2.5$ 6.74 ($\pm$ 1.70) 28.19% 58.58
ARGS + Lagrangian multiplier $\omega=2.5$ 3.21 ($\pm$ 1.59) 75.8% 63.42
ARGS + Cost Model $\omega=2.5$ 0.19 ($\pm$ 1.65) 81.6% 55.73
InferenceGuard 7.08 ($\pm$ 2.49) 88.14% 16.25
InferenceGuard with Critic 6.50 ($\pm$ 2.50) 94.46% 15.07
Beaver-7B-v3 Base 5.83 ($\pm$ 1.62) 75.89% 0.8
RE-Control 5.9 ($\pm$ 1.56) 75.9% 1.33
RE-Control + Lagrangian multiplier 5.91 ($\pm$ 1.50) 75.9% 1.73
Best-of-N + Lagrangian multiplier 6.58 ($\pm$ 1.89) 85.7% 14.5
Best-of-N + Augmented safety 9.05 ($\pm$ 1.59) 97.21% 9.62
Beam search + Lagrangian multiplier 8.63 ($\pm$ 1.80) 87.35 % 16.0
Beam search + Augmented safety 10.31 ($\pm$ 1.37) 97.36% 9.75
ARGS $\omega=2.5$ 6.72 ($\pm$ 1.83) 78.5% 67.15
ARGS $\omega=2.5$ + Lagrangian multiplier 2.26 ($\pm$ 1.56) 81% 72.58
ARGS $\omega=2.5$ + Cost Model 0.01 ($\pm$ 1.37) 98.4% 64.28
InferenceGuard 10.26 ($\pm$ 1.42) 99.7% 9.95
InferenceGuard with Ot Critic 10.27 ($\pm$ 1.50) 100% 9.76
Table 3: Performance Comparison using Vicuna-7B, evaluated by reward models llama-7b-rm and beaver-7b-unified-cost, on Datasets HEx-PHI and HH-RLHF $N=128$, $\lambda=5$